### v1

# Artificial Data Point Generation in Clustered Latent Space for Small Medical Datasets

Abstract
One of the growing trends in machine learning is the use of data generation techniques, since the performance of machine learning models is dependent on the quantity of the training dataset. However, in many medical applications, collecting large datasets is challenging due to resource constraints, which leads to overfitting and poor generalization. This paper introduces a novel method, Artificial Data Point Generation in Clustered Latent Space (AGCL), designed to enhance classification performance on small medical datasets through synthetic data generation. The AGCL framework involves feature extraction, K-means clustering, cluster evaluation based on a class separation metric, and the generation of synthetic data points from clusters with distinct class representations. This method was applied to Parkinson’s disease screening, utilizing facial expression data, and evaluated across multiple machine learning classifiers. Experimental results demonstrate that AGCL significantly improves classification accuracy compared to baseline, GN and kNNMTD. AGCL achieved the highest overall test accuracy of 83.33% and cross-validation accuracy of 90.90% in majority voting over different emotions, confirming its effectiveness in augmenting small datasets.
Index Terms:
Synthetic Data Generation, Data augmentation, Small datasets, Parkinson’s Disease Screening, K-means clustering, Normal Distribution
I Introduction
The performance of machine learning models heavily relies on the training dataset [1]. However, collecting a large amount of data is time-consuming and demands considerable resources. Unfortunately, many medical applications do not have enough data to properly train machine learning models. In many practical scenarios, where annotating labels extensively is labor-intensive or only limited datasets are accessible, machine learning models are prone to overfitting and poor generalization [2, 3].
In strict small dataset scenarios, the learner has access to a small number of labeled examples from each class, and the number of classes can also be quite small. In the fields of transfer learning, few-shot learning, and semi-supervised learning, several methods have been proposed. However, learning using small datasets is different from all three research areas. Methods designed to address strict small samples scenarios cannot rely on transfer learning from large amounts of peripheral data. The problem can be notably alleviated by imposing a strong prior on the model. Unfortunately, in many medical applications, we face an unknown domain where such prior knowledge is not available. In the few-shot scenario, the learner has access to a large number of labeled examples from classes not participating in the current classification task, while in the semi-supervised scenario, the learner typically has access to a large number of unlabeled examples [4, 5, 6].
Data augmentation and data generation techniques are effective strategies to address the issues of data scarcity [7]. Generating new synthetic samples is a valuable approach that enables models to generalize better and achieve higher accuracy when encountering unseen data. Leveraging generative adversarial networks to enhance a limited training dataset is particularly attractive, especially now that highly powerful deep generative models are increasingly available. However, these models typically demand a substantial sample size to train effectively, leading to poor performance in scenarios with limited data [8].
In this paper, we introduce the Artificial Data Point Generation in Clustered Latent Space (AGCL) as a solution to address the challenging problems of small medical datasets. The proposed method can be implemented as follows: first, perform feature extraction and then apply K-means clustering. We then assess the results of the clustering for class separation and, if necessary, perform re-clustering. Next, to introduce variety into the dataset, we create additional synthetic data points using the normal distribution based on each cluster’s parameters. Finally, the updated dataset, which combines original and synthetic data, is used to train machine learning models for improved classification performance.
The aim of this study is to tackle the issue of limited training data in the medical field, which makes it challenging to develop an effective classification model. While our primary focus in this paper is on Parkinson’s screening, our methodology can easily be applied to augmenting data in various other fields.
The rest of this article is structured as follows. Section II reviews related work. Section III outlines the methodology. Section IV describes the Dataset which we used for our experiments. Section V presents simulation results and discussions. Finally, Section VI concludes with key takeaways.
Figure 1: Illustrating the AGCL framework
II Related Work
When it comes to working with small datasets, data augmentation appears to be one of the few unmatched methods to boost the efficiency of machine learning models. There are many augmentation methods used for image classification, including random rotation, mirroring, and the addition of Gaussian noise [9]. Heuristic methods such as Cutout [10] partially occlude a square area in each training process to affect the learned features, and Random Erasing [11] partially covers or replaces certain areas of an image. Although these methods are very useful, they may not yield the best results on very small datasets and can negatively affect classification. Blending is another significant method of data augmentation that involves combining images or regions of interest within a single image. For example, the CutMix [12] technique replaces deleted pixels with portions of other images to create a good combination. However, problems can arise with methods such as CutMix, as the pasted patches can clash with the background of the original image and differ semantically from it, causing confusion during label assignment in the training process.
Automatic augmentation [13] has seen significant advancements, focusing on enhancing augmentation techniques through the use of reinforcement learning. This approach aims to automate the selection and fine-tuning of augmentation strategies, promising efficiency gains. However, a key challenge lies in the computational demands of the search algorithm to identify the optimal augmentation approach, particularly as data dimensions and model parameters increase. Feature augmentation represents another common approach to data enhancement, focusing on expanding learned features rather than the input space of the data. FitMatch [14] is one of the methods in this category which uses learned features to create different and elaborate transformations. Since FitMatch increases the number of features, this overloads the data dimensionality and the complexity of the models that would take more time to train when costs are incurred.
Addressing the problem of small datasets, the Least-Square Generative Adversarial Network (LS-GAN) was proposed. This algorithm begins with the independence test on feature removal by correlation measures and eliminates features that are linearly associated with other features. Next the feature selection method of Burato is used which employ random forest algorithm for the enhancement of feature selection. Yet another and rather important aspect of the generative network in the adversarial least squares sense is that the least squares cost function is utilized in this case, which do not allow vanishing gradient to come through [15].
Sivakumar et al. [16] have introduced a modified mega-trend diffusion to address the issues of generating artificial samples for small datasets. The algorithm involves identifying the k nearest neighbors for each sample in the dataset using the k-Nearest Neighbor and then applying Mega-Trend Diffusion to estimate the domain ranges of the neighboring samples. This is a multifunctional approach that can be used to provide artificial data needed in other data-centric tasks other than used in image classification tasks.
Although Oversampling [17] differs with the data generation, its algorithms can be employe in this area. Wei et al. [18] introduce the Improved and Random Synthetic Minority Oversampling Technique (IR-SMOTE). In data pre-processing, the K-means algorithm is utilized to cluster both majority and minority class samples, enabling the removal of noise samples in the minority class based on a distance metric. Subsequently, the kernel density estimation technique is employed to dynamically assign the number of synthetic samples to each cluster within the minority class. Finally, a synthetic approach that considers the attributes of the data is introduced to generate a more diverse set of synthetic samples.
III Methodology
As shown in fig 1, the general methodology of AGCL includes feature extraction, K-means clustering for grouping data samples, and assessment of the obtained clusters to identify class representation. The cluster evaluation is based on the separation criterion, with a threshold guiding the decision to further subdivide for improved class separation. These steps pave the way for the subsequent steps in the algorithm involved in data synthesis. The analysis of each part of the methodology is presented in the subsequent sections of the paper.
III-A Feature Extraction and Clustering
In the first step, relevant features are extracted from the dataset. Let $X=\{x_{1},x_{2},...,x_{n}\}$ be the dataset in the latent space. By K-means clustering, data points are organized into clusters $C=\{c_{1},c_{2},...,c_{k}\}$. The K-means algorithm operates through an iterative process aimed at minimizing the total sum of distances between each point in a cluster and its centroid [19]. By computing the mean of data points, new cluster centroids are determined, resulting in an arrangement where data points align effectively with their respective clusters. Each cluster $c_{j}$ comprises data points $x_{i,j}$ where $i=1,2,...,n_{j}$, with $n_{j}$ representing the number of data points in cluster $c_{j}$.
Moreover, all the hyperparameters of K-means clustering algorithm are also hyperparameters of the proposed method particularly $k$ the number of clusters. It is essential to select $k$ appropriately, ensuring it is greater than the number of classes present in the dataset. The determination of the number of clusters should initially be done using techniques like the elbow method.
III-B Cluster Evaluation and Re-Clustering Process
The next step involves identifying clusters containing points from a particular class. This evaluation aims to identify clusters that capture distinct classes within the dataset. When a cluster contains data points from only one class, it indicates that the cluster is homogeneous and well-separated from others, making it suitable for use in the generation step. However, if a cluster contains points from multiple classes, its quality needs to be evaluated further. For such mixed clusters, a separation criterion is applied to assess how well-separated the different classes are within the cluster. This helps decide whether the cluster should be further divided to improve its homogeneity. The separation criterion combines the Class Separation Metric (CSM) and entropy to provide a measure of cluster purity and class distribution.
The CSM for each cluster is defined as the ratio of inter-class separation to intra-class cohesion. Mathematically, it can be expressed as:
$\textit{CSM}(c_{j})=\frac{\mathcal{S}_{inter}(c_{j})}{\mathcal{C}_{intra}(c_{j})}$ (1)
where
$\mathcal{C}_{intra}(c_{j})=\sum\limits_{l\in\mathcal{L}}\frac{1}{n_{l}(n_{l}-1)}\sum\limits_{\begin{subarray}{c}i,k\in l\\ i\neq k\end{subarray}}\|x_{i}-x_{k}\|$ (2)
$\mathcal{S}_{inter}(c_{j})=\sum\limits_{\begin{subarray}{c}l_{i},l_{k}\in\mathcal{L}\\ l_{i}\neq l_{j}\end{subarray}}\frac{1}{n_{l_{i}}\cdot n_{l_{k}}}\sum\limits_{\begin{subarray}{c}x_{i}\in l_{i}\\ x_{k}\in l_{k}\end{subarray}}\|x_{i}-x_{k}\|$ (3)
The measure $\mathcal{C}_{intra}$ and $\mathcal{S}_{inter}$ are calculated separately for each cluster containing points from multiple classes. $\mathcal{C}_{intra}$ measures the average pairwise distance between all points with the same label within a cluster, where $n_{l}$ is the number of points in class $l$, and $\|x_{i}-x_{k}\|$ is the Euclidean distance between points within the same class. $\mathcal{S}_{inter}$ measures the average pairwise distance between points in different classes $l_{i}$ and $l_{k}$, where $n_{l_{i}}$ and $n_{l_{k}}$ are the numbers of points in classes $l_{i}$ and $l_{k}$, respectively.
A high CSM indicates that the clusters are well-separated in the feature space. This means that the average distance between data points from different classes is relatively large. This suggests that the classes are not heavily intermixed, and the data has some orders.
For each cluster, entropy can be calculated based on the distribution of class labels. A low entropy value indicates that most of the data points within the cluster belong to the same class, implying that the cluster is homogeneous. Conversely, a high entropy value suggests that the cluster contains a mix of different class labels, indicating that it is more heterogeneous or mixed.
For a given cluster $c_{j}$, the entropy $H(c_{j})$ can be defined as:
$H(c_{j})=-\sum\limits_{l\in\mathcal{L}}p(l)\log_{2}p(l)$ (4)
where $p(l)$ is the probability of a point belonging to class $l$ within the cluster. A low entropy value (close to 0) indicates that the cluster is mostly made up of data points from a single class, which is desirable for further dividing.
Finally, the separation criterion combines CSM and entropy metrics to assess the quality of a cluster in terms of both spatial separation and label distribution. It is defined as:
Separation $\textit{Separation Criterion}(c_{j})=\frac{CSM(c_{j})}{H(c_{j})}$ (5)
A high separation criterion indicates that the cluster is well-separated in the feature space. This means that the data points belonging to different classes are far apart from each other.
By normalizing it, the separation criterion ranges between zero and one. Selecting a suitable threshold for the separation criterion check is vital. In cases where a cluster does meet the specified threshold, a re-clustering process is triggered to improve the grouping of data points and enhance the separation of classes within that cluster. This iterative evaluation and re-clustering process aims to optimize the clustering results and identify better different class patterns. If it is observed that a specific cluster requires splitting into smaller clusters, the number of clusters for that particular cluster should be set to the number of classes in the dataset.
A threshold closer to one ensures fewer subdivisions in clustering, while a threshold closer to zero splits more existing clusters into smaller sub-clusters. For example, setting the threshold exactly at one would execute the K-means algorithm only once, while a threshold of zero would continue K-means clustering hierarchically until each cluster contains a single class.
This process is somewhat analogous to the bias-variance trade-off in machine learning. A threshold closer to one, resulting in fewer subdivisions, can be likened to higher bias. The model makes simpler assumptions about the data (fewer clusters), which might not capture all the nuances of the dataset. Conversely, a threshold closer to zero, leading to more subdivisions, can be compared to higher variance. The model makes more complex assumptions (more clusters), capturing more details and variations within the data, potentially leading to overfitting, as it may represent noise and subtle details too closely.
III-C Synthetic Data Generation
Clusters that contain data points from a single class are deemed suitable for generating synthetic data. On the other hand, clusters that contain data points from multiple classes after assessing separation criterion are considered mixed and are typically discarded. We produce synthetic data points based on the estimated parameters of clusters containing points from a particular class using normal distribution. It depicts the process of obtaining new artificial data points starting with clustering radius and the center to fit the observed distribution.
For clusters with more than one data point, the cluster radius ($r_{j}$) is calculated based on the maximum distance between any point in the cluster and the cluster center ($\mu_{j}$), plus a small augmentation factor proportional to the range of these distances. This ensures that the radius takes into account the spread of points around the center, while also allowing for a slight margin beyond the maximum distance. For clusters with only a single data point, the radius is set to a constant value, which is derived from the average standard deviation of the features, scaled by a small factor. Therefore, the radius $r_{j}$ for each cluster $j$ is defined as:
:$:\left\{r_{j}\leftarrow\begin{cases}0.01\times\frac{1}{d}\sum_{i=1}^{d}\sigma_{j}&\text{if }n_{c_{j}}==1\\ \max(D_{x,\mu_{j}})+0.1\times(\max(D_{x,\mu_{j}})-\min(D_{x,\mu_{j}}))&\text{if }n_{c_{j}}>1\end{cases}\\ \right.$
$\sigma_{j}$ is the standard deviation of the $j$-th feature calculated based on the data points from the previous clustering step. This indicates that the variability is measured considering the feature distribution observed in the last iteration or phase of clustering. Also, $D_{x,\mu_{j}}$ refers to the distance of each point $\begin{subarray}{c}x\in X_{j}\end{subarray}$ from the cluster center $\mu_{j}$.
The generation of synthetic data points involves sampling $q$ artificial data points from a normal distribution centered at the cluster mean ($\mu_{j}$), as depicted in Equation 7:
$x_{new,q}\sim N(\mu_{j},\Sigma_{j})$ (7)
$\Sigma_{j}$ represents the covariance matrix calculated based on the variance of $r_{j}$.
The number of synthetic data points generated in each cluster is proportional to $n_{c_{j}}$, denoted by the coefficient $\alpha$.
$q=\alpha\cdot n_{c_{j}}$ (8)
The coefficient $\alpha$ controls the number of synthetic data points against the number of data points of the given set. It is flexible according to requirements of a specific dataset and clustering needs. This proportional augmentation approach makes it possible to assign more synthetic data points to areas of the input space that are densely occupied by training data points. By applying this customized augmentation strategy, the clustering algorithm is better equipped to handle outliers and variations in cluster sizes.
III-D Dataset Update and Classification
The dataset is updated with the synthetic data points to form $X_{augmented}\leftarrow X\cup\{x_{synth}\}$, providing an enriched set of training data for machine learning models. Leveraging machine learning models on the updated dataset enables the classification tasks for identifying patterns and making predictions.
In summary, AGCL tackles the challenges of small medical datasets by combining feature extraction, iterative K-means clustering, and synthetic data generation. The augmented dataset, comprising both original and synthetic data points, is utilized to train machine learning models, enhancing their performance in classification tasks. Ultimately, the algorithm is presented generically in Algorithm 1.
Input : Dataset $X$, Labels $y$, Parameters $\textit{thrsh},\alpha$
Output : Augmented dataset $X_{augmented}$
$X\leftarrow\text{FeatureExtractor}(X)$
Fit K-means on $X$ to obtain Clusters $C=\{c_{1},c_{2},...,c_{k}\}$ with centers $\{\mu_{1},\mu_{2},...,\mu_{k}\}$
$X_{synth}\leftarrow\emptyset$
for $j=1$ to $k$ do
       $X_{j}\leftarrow\{x\mid x\in X\text{ and }x\in c_{j}\}$
       $y_{j}\leftarrow\{y\mid x\in c_{j}\}$
       $n_{c_{j}}\leftarrow\text{size}(X_{j})$
      if $\text{Unique}(y_{j})$ is False  then
             Separation $\textit{Separation Criterion}\leftarrow\frac{\text{CSM}(c_{j})}{\text{H}(c_{j})}$
[…]
                   Call the algorithm recursively with $X_{j}$, $y_{j}$
      else
$r_{j}\leftarrow\begin{cases}0.01\times\frac{1}{d}\sum_{i=1}^{d}\sigma_{j}&\text{if }n_{c_{j}}=1\\ \max(D_{x,\mu_{j}})+0.1\times(\max(D_{x,\mu_{j}})-\min(D_{x,\mu_{j}}))&\text{if }n_{c_{j}}>1\end{cases}$
             $q\leftarrow\alpha\cdot n_{c_{j}}$
             $\Sigma_{j}\leftarrow\text{diag}(r_{j}^{2})$
            for $i=1$ to $q$ do
                   $x_{new}\sim N(\mu_{j},\Sigma_{j})$
                   $X_{synth}\leftarrow X_{synth}\cup\{x_{new}\}$
$X_{augmented}\leftarrow X\cup X_{synth}$
Algorithm 1 AGCL Pseudocode
IV Dataset
Parkinson’s disease is a prevalent clinical syndrome that belongs to the group of neurodegenerative movement disorders, manifesting as progressive motor disability in the elderly. This condition is characterized by the progressive decay of nerve cells in the substantia nigra, a part of the brain, reducing the manufacture of dopamine, a chemical that plays an important role in the control of body movements [20, 21, 22].
Parkinson’s disease significantly impacts the motor system, with key symptoms including bradykinesia, rest tremor, stiffness, and rigidity that often appear in the early stages of the condition. Bradykinesia is characterized by reduced and slowed movements, which can also affect the facial muscles in PD patients, making it challenging to produce facial expressions. This reduction in facial expression is commonly referred to as hypomimia or masked faces [23, 24].
The Facial Action Coding System (FACS) is a detailed and anatomically based framework designed to categorize and analyze all observable facial movements. It deconstructs facial expressions into specific components of muscle movement known as Action Units (AUs) [25]. Researchers have utilized changes in Action Units associated with hypomimia to effectively screen Parkinson’s disease and identify hypomimia in affected individuals [26].
PD screening can be aided by analyzing videos and photos capturing the individual’s facial expressions while showing specific emotions. Usually, individuals are asked to express basic emotions like happiness, sadness, or disgust. By studying these images, conditions such as hypomimia can be detected.
Our dataset consists of videos from 55 individuals captured using an Android tablet camera. Prior to the task, participants viewed a brief video of a person performing the task. They were then asked to make a facial expression and hold it for a few seconds. The facial expressions include anger, happiness, disgust, fear, and surprise. Examples of these facial expressions are illustrated in Figure 2. Not all participants completed all five facial expressions; therefore, the distribution of videos for each emotion is presented in Table I. Since each video represents a transition from a neutral face to an extreme emotion, the peak frame of each video is selected for analysis.
Figure 2: Facial expressions from left to right: anger, happy, disgust, fear and surprise
TABLE I: Dataset Characteristics
Emotion Rage Happy Disgust Fear Surprise
Parkinson’s (PD) 26 25 26 25 29
Control 28 30 27 30 26
TABLE II: Associated AUs to each facial expression
Emotion Facial Action Unit
Rage AU09 (Nose Wrinkler)
AU10 (Upper Lip Raiser)
AU20 (Lip Stretcher)
Laugh AU15 (Lip Corner Depressor)
AU17 (Chin Raiser)
AU25 (Lips Part)
Disgust AU23 (Lip Tightener)
AU25 (Lips Part)
AU45 (Blink)
Fear AU05 (Upper Lid Raiser)
AU12 (Lip Corner Puller)
AU17 (Chin Raiser)
Surprise AU04 (Brow Lowerer)
AU05 (Upper Lid Raiser)
AU15 (Lip Corner Depressor)
V Experiments
In this section, we evaluate the AGCL method in various scenarios. Our objective is to provide a thorough assessment of AGCL’s performance under diverse conditions.
V-A Feature Extraction and Latent Space
Utilizing Openface [27], we extracted action unit (AU) values for each participant’s picture. Facial action units are linked to specific muscle movements of the face, with each unit corresponding to the movement of a distinct group of facial muscles. For example, activation of AU1 (also known as Outer Brow Raiser) indicates the simultaneous movement of two facial muscles - the frontalis and pars lateralis. The OpenFace software provides a binary activation (0 or 1) and a raw magnitude (ranging from 0 to 5) for each AU for every frame of a video containing a human face.
A facial expression can be associated with multiple action units. In this paper, each facial expression is associated with three Aus by analyzing Aus’ mean, concept of each AU associated with facial expression and Pearson Correlation. Table II shows the AUs associated with facial expressions.
TABLE III: Comparison between features of real and synthetic data based on t-test, Levene-test and ks-test
t-test (p-value) Levene-test (p-value) ks-test (p-value)
Emotion Feature Control Parkinson Control Parkinson Control Parkinson
Rage AU09 0.77 0.72 0.71 0.71 0.33 0.83
AU10 0.83 0.10 0.97 0.78 0.83 0.17
AU20 0.36 0.79 0.71 0.34 0.57 0.57
Happy AU15 0.21 0.64 0.23 0.72 0.01 0.98
AU17 0.01 0.53 0.01 0.94 0.01 0.57
AU25 0.22 0.82 0.47 0.32 0.33 0.83
Disgust AU23 0.36 0.48 0.32 0.54 0.33 0.83
AU25 0.14 0.30 0.08 0.83 0.17 0.33
AU45 0.07 0.99 0.13 0.46 0.57 0.98
Fear AU05 0.87 0.27 0.85 0.05 0.98 0.33
AU12 0.46 0.72 0.45 0.24 0.17 0.57
AU17 0.31 0.16 0.31 0.52 0.98 0.57
Surprise AU04 0.92 0.35 0.92 0.39 0.33 0.83
AU05 0.19 0.70 0.54 0.65 0.33 0.83
AU15 0.85 0.15 0.26 0.77 0.17 0.33
V-B Statistical evaluations
Three different tests were conducted to verify that the real data and synthetic data do not differ statistically. Twenty random samples were selected from each class of the two datasets (real and synthetic). The p-values for each test are shown in the table III. In the t-test [28], it is observed that the p-value is above 0.05 for all variables except one, indicating that the means of the two datasets are not significantly different. Levene’s test [29] was used to compare the variances between the datasets. The p-value is above 0.05 for all variables except one, suggesting that the variances are not statistically significant. Similarly, the most p-values from the Kolmogorov-Smirnov test (ks-test) [30] are also above 0.05, showing that the distribution of variables in the two datasets is not statistically significant. Therefore, we can conclude that the two datasets are statistically similar.
V-C Classification
TABLE IV: Comparison between accuracy of AGCL, GN, kNNMTD and Baseline on 5 different emotions.
KNN SVM RF MLP
Emotion Method Validation Test Validation Test Validation Test Validation Test
Rage Baseline 54.00 ($\pm$39.80) 69.23 52.00 ($\pm$38.68) 53.84 64.00 ($\pm$38.78) 46.15 66.00 ($\pm$33.82) 53.84
GN 58.00 ($\pm$39.19) 69.23 56.00 ($\pm$38.26) 53.84 68.00 ($\pm$37.09) 46.15 72.00 ($\pm$31.87) 46.15
kNNMTD 54.00 ($\pm$39.80) 69.23 56.00 ($\pm$38.26) 53.84 64.00 ($\pm$38.78) 53.84 54.00 ($\pm$42.24) 61.53
AGCL 60.00 ($\pm$40.00) 76.92 58.00 ($\pm$41.67) 69.23 68.00 ($\pm$39.70) 53.84 74.00 ($\pm$32.00) 53.84
Happy Baseline 70.83 ($\pm$40.61) 61.11 70.83 ($\pm$43.10) 72.22 79.17 ($\pm$35.11) 66.66 75.00 ($\pm$38.19) 66.67
GN 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$35.11) 72.22 79.17 ($\pm$35.11) 66.66 81.25 ($\pm$31.66) 77.78
kNNMTD 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$37.96) 72.22 81.25 ($\pm$34.80) 66.68 68.75 ($\pm$42.85) 61.11
AGCL 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$35.11) 83.33 81.25 ($\pm$34.80) 72.22 81.25 ($\pm$31.66) 83.33
Disgust Baseline 68.00 ($\pm$37.09) 46.15 76.00 ($\pm$32.00) 46.15 78.00 ($\pm$28.57) 61.54 70.00 ($\pm$40.00) 69.23
GN 76.00 ($\pm$34.99) 53.85 80.00 ($\pm$31.62) 38.46 82.00 ($\pm$24.00) 69.23 82.00 ($\pm$27.86) 30.77
kNNMTD 68.00 ($\pm$37.09) 46.15 76.00 ($\pm$34.99) 46.15 84.00 ($\pm$23.32) 61.54 72.00 ($\pm$37.63) 61.54
AGCL 72.00 ($\pm$37.63) 69.23 82.00 ($\pm$31.24) 53.84 84.00 ($\pm$30.72) 69.23 84.00 ($\pm$27.28) 76.92
Fear Baseline 60.00 ($\pm$40.00) 69.23 68.00 ($\pm$37.09) 84.62 62.00 ($\pm$38.16) 61.53 62.00 ($\pm$38.16) 53.84
GN 64.00 ($\pm$36.11) 69.23 66.00 ($\pm$39.29) 76.92 66.00 ($\pm$36.66) 61.53 70.00 ($\pm$37.42) 76.92
kNNMTD 60.00 ($\pm$40.00) 69.23 68.00 ($\pm$39.70) 84.62 72.00 ($\pm$31.87) 53.84 68.00 ($\pm$39.70) 61.53
AGCL 62.00 ($\pm$38.16) 76.92 70.00 ($\pm$40.00) 84.62 74.00 ($\pm$37.74) 69.23 72.00 ($\pm$37.63) 92.30
Surprise Baseline 68.00 ($\pm$39.70) 56.25 78.00 ($\pm$34.87) 43.75 70.00 ($\pm$34.64) 43.75 72.00 ($\pm$34.87) 62.50
GN 72.00 ($\pm$37.63) 50.00 70.00 ($\pm$37.42) 43.75 80.00 ($\pm$31.62) 50.00 76.00 ($\pm$34.99) 43.75
kNNMTD 68.00 ($\pm$39.70) 56.25 68.00 ($\pm$39.70) 50.00 68.00 ($\pm$34.29) 50.00 72.00 ($\pm$37.63) 56.25
AGCL 70.00 ($\pm$37.42) 62.50 78.00 ($\pm$34.87) 50.00 82.00 ($\pm$31.24) 56.25 84.00 ($\pm$30.72) 75.00
Voting Baseline 68.18 61.11 75.00 66.66 77.27 61.11 65.90 72.22
GN 70.45 66.66 70.45 61.11 81.81 66.66 77.27 66.66
kNNMTD 68.18 61.11 77.27 66.66 79.54 66.66 77.27 66.66
AGCL 75.00 77.77 79.54 77.77 90.90 77.77 84.09 83.33
In the previous section, we analyzed the data statistically. In this section, we evaluate how the data behaves when subjected to machine learning algorithms for classifying Parkinson’s and control data. A two-leave-out cross-validation method was applied to the training data, resulting in 25 unique folds. Each fold served as an independent subset for training separate models using the same hyperparameters. Throughout the cross-validation process, hyperparameter tuning was carried out to identify the optimal hyperparameters based on performance across all folds. The test data consisted of 11 control samples and 7 Parkinson’s disease samples. Instead of selecting a single model, all 25 models generated during the cross-validation process were applied to the test data. Their predictions were aggregated using a weighted majority voting approach to produce the final classification.
By comparing the effectiveness of AGCL against baseline, perturbation mechanisms and kNNMTD [16], we aim to evaluate its performance in classifying facial expressions. The baseline approach serves as a reference point without any augmentation, providing a foundational measure of classification performance.
In contrast, the perturbation-based synthetic data generation mechanism is stated as: $x^{s}=x_{i}+\epsilon$ where $\epsilon$ is a noise sampled from a certain distribution. According to Gaussian approach, $\epsilon$ is sampled with $N(0,\sigma)$. This method usually utilizes n-way marginals for purposes of creating variability in the categorical features [31]. Therefore, we employed the Gaussian approach to add noise to each sample in the latent space. To elucidate the impact of different levels of noise, we consider a range of standard deviations ($\sigma$= 0.001, 0.005, 0.01, 0.025, 0.05, 0.075, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.8, 1]) to identify the optimal performance of Gaussian Noise (GN) for comparison with the AGCL method.
Moreover, We estimated the number of generated data points in the cluster for AGCL through grid search, gradually testing the $\alpha$ value from 2 to 60 with a step size of 5.
Table IV shows the performance of algorithms across four machine learning classifiers in screening Parkinson across five emotions: Rage, Happy, Disgust, Fear, and Surprise. To aggregate results obtained from five facial expressions, we employ the majority voting integration method, by voting from all facial expressions for each individual. This ensemble method helps in reducing the possibility of bias or errors that come with separate facial expressions, making it a more accurate and diverse approach. The metrics reported include both cross-validation and test accuracy.
For the rage emotion, AGCL demonstrated superior performance compared to other methods across all classifiers. The best validation accuracy was achieved using the MLP classifier, where AGCL reached 74.00%, whereas the baseline validation accuracy was lower at 66.00%. In the test results, AGCL also led with the highest accuracy of 76.92% with kNN, outperforming the other algorithms test accuracy of 69.23%.
In the happy emotion, AGCL achieved the highest validation accuracy of 81.25% with MLP, significantly improving over the kNNMTD’s validation accuracy of 68.75%. The test accuracy of 83.33% with both SVM and MLP shows the effectiveness of AGCL in generating high-quality synthetic data that enhances classification performance, compared to the GN and kNNMTD. While AGCL performed well in test accuracy compared to the baseline and other augmentation methods, the overall improvement in cross-validation was less pronounced than in other emotions. The validation accuracy of 81.25% with RF and MLP showed moderate gains, which suggests that the happy expression might be easier for classifiers to distinguish even without extensive data augmentation.
For the disgust emotion, AGCL also led to noticeable performance improvements. The best validation accuracy of 84.00% was achieved using the Random Forest classifier. AGCL also reached a test accuracy of 76.92% with MLP, which is the best test accuracy across all four algorithms. The fear emotion further highlights the advantages of AGCL in small dataset classification. AGCL achieved the highest validation accuracy of 74.00% with the Random Forest classifier, while the GN struggled with only 66.00%. On the test set, AGCL achieved its best accuracy with MLP, reaching an outstanding 92.30%, far exceeding the other algorithms.
In the surprise emotion, AGCL continued to outperform the baseline and other methods. The highest validation accuracy was 84.00% with MLP for AGCL. In the test phase, AGCL again demonstrated strong performance, achieving 75.00% accuracy with MLP, surpassing the baseline’s test accuracy of 62.50%. When results across all emotions were aggregated using majority voting, AGCL achieved the highest overall test accuracy of 83.33% with the MLP classifier. This marks a significant improvement over the other algorithms. The validation accuracy of 90.90% with Random Forest and 84.09% with MLP further confirms AGCL’s effectiveness in augmenting small medical datasets and improving classification results.
VI Conclusion
In this paper, we introduced the Artificial Data Point Generation in Clustered Latent Space (AGCL) method to address the challenges of limited training data in medical applications, particularly in the classification of Parkinson’s disease through facial expression analysis. By generating synthetic data points in a clustered latent space, AGCL enhances the performance of machine learning models in scenarios with small datasets. The framework employs feature extraction, K-means clustering, and a separation criterion to guide synthetic data generation, ensuring that only well-separated clusters contribute to data augmentation.
Experimental results across multiple classifiers and emotions demonstrate that AGCL consistently outperforms baseline methods, as well as other data augmentation techniques like Gaussian noise and kNNMTD. AGCL achieved the highest overall test accuracy of 83.33% through majority voting, confirming its effectiveness in improving classification in small datasets.
AGCL’s ability to generate synthetic data points tailored to the structure of the dataset offers significant advantages in enhancing the generalization capabilities of machine learning models. This method holds promise for broader applications in medical domains and other fields where data scarcity is a critical issue. Future work may explore the extension of AGCL to different datasets and further refinement of clustering techniques to increase its applicability across a wider range of tasks.

### v2

@@
-Artificial Data Point Generation in Clustered Latent Space for Small Medical Datasets
+Feature-to-Image Data Augmentation: Improving Model Feature Extraction with Cluster-Guided Synthetic Samples
 Abstract
 One of the growing trends in machine learning is the use of data generation techniques, since the performance of machine learning models is dependent on the quantity of the training dataset.
-However, in many medical applications, collecting large datasets is challenging due to resource constraints, which leads to overfitting and poor generalization.
-This paper introduces a novel method, Artificial Data Point Generation in Clustered Latent Space (AGCL), designed to enhance classification performance on small medical datasets through synthetic data generation.
-The AGCL framework involves feature extraction, K-means clustering, cluster evaluation based on a class separation metric, and the generation of synthetic data points from clusters with distinct class representations.
-This method was applied to Parkinson’s disease screening, utilizing facial expression data, and evaluated across multiple machine learning classifiers.
-Experimental results demonstrate that AGCL significantly improves classification accuracy compared to baseline, GN and kNNMTD.
-AGCL achieved the highest overall test accuracy of 83.33% and cross-validation accuracy of 90.90% in majority voting over different emotions, confirming its effectiveness in augmenting small datasets.
+However, in many real-world applications, particularly in medical and low-resource domains, collecting large datasets is challenging due to resource constraints, which leads to overfitting and poor generalization.
+This study introduces FICAug, a novel feature-to-image data augmentation framework designed to improve model generalization under limited data conditions by generating structured synthetic samples.
+FICAug first operates in the feature space, where original data are clustered using the k-means algorithm.
+Within pure-label clusters, synthetic data are generated through Gaussian sampling to increase diversity while maintaining label consistency.
+These synthetic features are then projected back into the image domain using a generative neural network, and a convolutional neural network is trained on the reconstructed images to learn enhanced representations.
+Experimental results demonstrate that FICAug significantly improves classification accuracy.
+In feature space, it achieved a cross-validation accuracy of 84.09%, while training a ResNet-18 model on the reconstructed images further boosted performance to 88.63%, illustrating the effectiveness of the proposed framework in extracting new and task-relevant features.
 Index Terms:
-Synthetic Data Generation, Data augmentation, Small datasets, Parkinson’s Disease Screening, K-means clustering, Normal Distribution
+Data augmentation, Small datasets, Parkinson’s Disease Screening, Self-Supervised Learning, Ganimation
 I Introduction
-The performance of machine learning models heavily relies on the training dataset [1].
-However, collecting a large amount of data is time-consuming and demands considerable resources.
-Unfortunately, many medical applications do not have enough data to properly train machine learning models.
-In many practical scenarios, where annotating labels extensively is labor-intensive or only limited datasets are accessible, machine learning models are prone to overfitting and poor generalization [2, 3].
-In strict small dataset scenarios, the learner has access to a small number of labeled examples from each class, and the number of classes can also be quite small.
-In the fields of transfer learning, few-shot learning, and semi-supervised learning, several methods have been proposed.
-However, learning using small datasets is different from all three research areas.
-Methods designed to address strict small samples scenarios cannot rely on transfer learning from large amounts of peripheral data.
-The problem can be notably alleviated by imposing a strong prior on the model.
-Unfortunately, in many medical applications, we face an unknown domain where such prior knowledge is not available.
-In the few-shot scenario, the learner has access to a large number of labeled examples from classes not participating in the current classification task, while in the semi-supervised scenario, the learner typically has access to a large number of unlabeled examples [4, 5, 6].
-Data augmentation and data generation techniques are effective strategies to address the issues of data scarcity [7].
-Generating new synthetic samples is a valuable approach that enables models to generalize better and achieve higher accuracy when encountering unseen data.
-Leveraging generative adversarial networks to enhance a limited training dataset is particularly attractive, especially now that highly powerful deep generative models are increasingly available.
-However, these models typically demand a substantial sample size to train effectively, leading to poor performance in scenarios with limited data [8].
-In this paper, we introduce the Artificial Data Point Generation in Clustered Latent Space (AGCL) as a solution to address the challenging problems of small medical datasets.
-The proposed method can be implemented as follows: first, perform feature extraction and then apply K-means clustering.
-We then assess the results of the clustering for class separation and, if necessary, perform re-clustering.
-Next, to introduce variety into the dataset, we create additional synthetic data points using the normal distribution based on each cluster’s parameters.
-Finally, the updated dataset, which combines original and synthetic data, is used to train machine learning models for improved classification performance.
-The aim of this study is to tackle the issue of limited training data in the medical field, which makes it challenging to develop an effective classification model.
-While our primary focus in this paper is on Parkinson’s screening, our methodology can easily be applied to augmenting data in various other fields.
-The rest of this article is structured as follows.
-Section II reviews related work.
-Section III outlines the methodology.
-Section IV describes the Dataset which we used for our experiments.
-Section V presents simulation results and discussions.
-Finally, Section VI concludes with key takeaways.
-Figure 1: Illustrating the AGCL framework
+One of the significant challenges in machine learning is collecting high-quality, diverse, and labeled data.
+The performance of most models relies directly on the quality and quantity of training data [1].
+In many real-world applications—especially in sensitive domains like healthcare—data are scarce, difficult to acquire, and prone to have legal, ethical, and budgetary constraints.
+The challenge becomes even more severe when dealing with rare diseases or specific populations, few labeled examples are available.
+These limitations highlight the importance of data-efficient methods that can handle well even when only a little training data is available [2, 3, 4].
+To address data scarcity, various augmentation and synthetic data generation techniques have been proposed.
+Among them, generating new data in the feature space is a promising strategy [5].
+In feature space, each data sample is mapped to a set of descriptive features, and new data points can be synthetically created by leveraging distributions that capture intra-class variability.
+Feature space generation provides a controlled and label-consistent sampling, making this methodology very useful for enlarging small training sets without disrupting the underlying structure of the data.
+In this study, we introduce the FICAug (feature-to-image clustered augmentation) pipeline designed to improve classification performance on small datasets, with a focus on Parkinson’s disease screening.
+The FICAug pipneline begins by clustering the original feature space using the k-means algorithm, followed by evaluation and refinement to ensure cluster purity.
+Within each clusters that comes as single label, synthetic feature vectors are generated using gaussian sampling.
+The synthetic features are mapped from the feature-space to image domain through a generative neural network in order to push the model learn new and more discriminative features that are not present in our original feature representation.
+This pipeline enables training a convolutional neural network (CNN) on the reconstructed images, leading to improved performance and generalization.
+By using this approach, the model learns new and informative representations from the structured synthetic data derived from existing features.
+Our main objective is to demonstrate that sample from latent meaningful clusters and learning with the reconstruction from other data will enhance the performance of models.
+Experiments show that the proposed method improves classification performance and offers a novel solution to the challenge of learning from limited data.
+This paper is focused on Parkinson’s disease screening but the proposed approach is general enough to be applied in different domains with small data.
+The rest of this article is organized as follows.
+In Section II, we provide an overview of existing research on data augmentation.
+Section III details our proposed method, including the clustering-based feature augmentation process and the generative reconstruction strategy used to produce image-space data.
+Section IV introduces the dataset used in our experiments, along with its characteristics and preprocessing steps.
+Section V presents the experimental setup and performance analysis.
+Finally, Section VI summarizes the key findings.
 II Related Work
-When it comes to working with small datasets, data augmentation appears to be one of the few unmatched methods to boost the efficiency of machine learning models.
+Small dataset learning is different from the closely related task of transfer learning, few-shot learning and semi-supervised methods.
+Unlike transfer learning, where models benefit from large-scale pretraining, or few-shot learning, where auxiliary classes with many labeled samples support generalization, small-data learning typically lacks access to both peripheral data and domain-specific priors.
+Semi-supervised learning, on the other hand, assumes the availability of a large pool of unlabeled data—which is often not the case in sensitive applications like medicine [6, 7, 8].
+When it comes to working with small datasets, data augmentation appears to be one of the typical methods to improve the performance of machine learning models.
 There are many augmentation methods used for image classification, including random rotation, mirroring, and the addition of Gaussian noise [9].
 Heuristic methods such as Cutout [10] partially occlude a square area in each training process to affect the learned features, and Random Erasing [11] partially covers or replaces certain areas of an image.
-Although these methods are very useful, they may not yield the best results on very small datasets and can negatively affect classification.
-Blending is another significant method of data augmentation that involves combining images or regions of interest within a single image.
-For example, the CutMix [12] technique replaces deleted pixels with portions of other images to create a good combination.
-However, problems can arise with methods such as CutMix, as the pasted patches can clash with the background of the original image and differ semantically from it, causing confusion during label assignment in the training process.
-Automatic augmentation [13] has seen significant advancements, focusing on enhancing augmentation techniques through the use of reinforcement learning.
-This approach aims to automate the selection and fine-tuning of augmentation strategies, promising efficiency gains.
-However, a key challenge lies in the computational demands of the search algorithm to identify the optimal augmentation approach, particularly as data dimensions and model parameters increase.
-Feature augmentation represents another common approach to data enhancement, focusing on expanding learned features rather than the input space of the data.
-FitMatch [14] is one of the methods in this category which uses learned features to create different and elaborate transformations.
-Since FitMatch increases the number of features, this overloads the data dimensionality and the complexity of the models that would take more time to train when costs are incurred.
-Addressing the problem of small datasets, the Least-Square Generative Adversarial Network (LS-GAN) was proposed.
-This algorithm begins with the independence test on feature removal by correlation measures and eliminates features that are linearly associated with other features.
-Next the feature selection method of Burato is used which employ random forest algorithm for the enhancement of feature selection.
-Yet another and rather important aspect of the generative network in the adversarial least squares sense is that the least squares cost function is utilized in this case, which do not allow vanishing gradient to come through [15].
-Sivakumar et al.
-[16] have introduced a modified mega-trend diffusion to address the issues of generating artificial samples for small datasets.
-The algorithm involves identifying the k nearest neighbors for each sample in the dataset using the k-Nearest Neighbor and then applying Mega-Trend Diffusion to estimate the domain ranges of the neighboring samples.
-This is a multifunctional approach that can be used to provide artificial data needed in other data-centric tasks other than used in image classification tasks.
-Although Oversampling [17] differs with the data generation, its algorithms can be employe in this area.
-Wei et al.
-[18] introduce the Improved and Random Synthetic Minority Oversampling Technique (IR-SMOTE).
-In data pre-processing, the K-means algorithm is utilized to cluster both majority and minority class samples, enabling the removal of noise samples in the minority class based on a distance metric.
-Subsequently, the kernel density estimation technique is employed to dynamically assign the number of synthetic samples to each cluster within the minority class.
-Finally, a synthetic approach that considers the attributes of the data is introduced to generate a more diverse set of synthetic samples.
+Mixing is another significant method of data augmentation that involves combining images or regions of interest within a single image.
+These methods typically operate by combining two or more images—or subregions of them—into a single composite image.
+One widely used approach is Mixup [12], which generates weighted combinations of random image pairs by linearly interpolating both their pixel values and labels.
+Another method, known as CutMix [13], replaces a randomly selected region of an image with a patch from another image.
+Recent advancements in automatic augmentation aim to optimize augmentation strategies using reinforcement learning, offering potential efficiency improvements [14].
+However, these methods often face high computational costs due to the complexity of searching for optimal policies.
+Alternatively, feature augmentation techniques such as FitMatch [15] generate new representations directly in feature space.
+While effective, they may significantly increase data dimensionality and model complexity, leading to longer training times.
+Generative adversarial networks (GANs) offer a promising solution for augmenting datasets, especially with the growing availability of powerful generative models [16].
+However, their effectiveness is often limited in low-data settings due to the large sample size typically required for training.
+To address this, the Least-Square GAN (LS-GAN) was introduced.
+It starts with feature decorrelation using statistical tests, followed by Burato’s feature selection method based on random forests.
+LS-GAN also employs a least-squares loss to prevent gradient vanishing and improve training stability [17].
+Feature space data augmentation has become an effective strategy for small datasets because it offers more control over data distribution than raw input-space augmentation.
+A simple approach in this domain is perturbation, where techniques like Gaussian noise injection simulate variability in a controlled way [18].
+For more structure, the kNNMTD method [19] generates synthetic data using $k$-nearest neighbors and Mega-Trend Diffusion.
+It selects neighbor points and uses membership functions to ensure synthetic samples preserve the original data relationships.
+Moreover, MEML [20] addresses deep learning contexts by generating synthetic features within intermediate network layers.
+It computes class-wise feature means and extrapolates between them and actual samples to produce meaningful representations during training.
+Autoencoders are also used for synthetic data generation.
+DeVries and Taylor [21] apply transformations like Gaussian noise, interpolation, and extrapolation to learned context vectors.
+This allows the creation of diverse synthetic samples while maintaining semantic alignment with the original data.
+Self-supervised learning methods have emerged as powerful tools for extracting rich representations without relying on labeled data.
+Two prominent examples are DINO [22] and Masked siamese network(MSN) [23].
+DINO leverages Vision Transformers in a self-distillation framework, encouraging different views of the same image to produce similar embeddings through teacher–student networks.
+This allows the model to capture diverse and meaningful features applicable to downstream tasks without supervision.
+Similarly, MSN proposes a Masked Siamese Network that learns representations by comparing masked views of images through a contrastive-style architecture, emphasizing label-efficient learning.
+Both DINO and MSN are self-supervised learning frameworks designed to extract general-purpose features from large-scale unlabeled datasets.
+These features are often used in downstream tasks with or without fine-tuning.
 III Methodology
-As shown in fig 1, the general methodology of AGCL includes feature extraction, K-means clustering for grouping data samples, and assessment of the obtained clusters to identify class representation.
-The cluster evaluation is based on the separation criterion, with a threshold guiding the decision to further subdivide for improved class separation.
-These steps pave the way for the subsequent steps in the algorithm involved in data synthesis.
-The analysis of each part of the methodology is presented in the subsequent sections of the paper.
+Figure 1: Overview of the FICAug framework, which generates synthetic feature vectors through clustering and Gaussian sampling, reconstructs them into images via a generative model, and trains a CNN to extract enhanced task-specific features.
+As shown in Figure 1, the general methodology of FICAug includes generating synthetic data points in the feature space by applying k-means clustering to the original dataset and identifying label-consistent regions.
+Within these clusters, new samples are created using Gaussian sampling to enhance intra-class diversity while maintaining the underlying distribution.
+These synthetic features are then projected back into the image domain using a generative neural network.
+A convolutional neural network (CNN) is trained on these reconstructed images, enabling the model to extract richer and more task-specific features from structured synthetic data.
 III-A Feature Extraction and Clustering
-In the first step, relevant features are extracted from the dataset.
+In the first step, the relevant features are extracted from the dataset.
 Let $X=\{x_{1},x_{2},...,x_{n}\}$ be the dataset in the latent space.
@@
 The model makes more complex assumptions (more clusters), capturing more details and variations within the data, potentially leading to overfitting, as it may represent noise and subtle details too closely.
-III-C Synthetic Data Generation
+III-C Synthetic Data Generation in Feature Space
 Clusters that contain data points from a single class are deemed suitable for generating synthetic data.
@@
 Therefore, the radius $r_{j}$ for each cluster $j$ is defined as:
-:$:\left\{r_{j}\leftarrow\begin{cases}0.01\times\frac{1}{d}\sum_{i=1}^{d}\sigma_{j}&\text{if }n_{c_{j}}==1\\ \max(D_{x,\mu_{j}})+0.1\times(\max(D_{x,\mu_{j}})-\min(D_{x,\mu_{j}}))&\text{if }n_{c_{j}}>1\end{cases}\\ \right.$
+$r_{j}\leftarrow\begin{cases}0.01\times\frac{1}{d}\sum_{i=1}^{d}\sigma_{j}&\text{if }n_{c_{j}}==1\\ \max(D_{x,\mu_{j}})+0.1\times(\max(D_{x,\mu_{j}})-\min(D_{x,\mu_{j}}))&\text{if }n_{c_{j}}>1\end{cases}\$
 $\sigma_{j}$ is the standard deviation of the $j$-th feature calculated based on the data points from the previous clustering step.
@@
 By applying this customized augmentation strategy, the clustering algorithm is better equipped to handle outliers and variations in cluster sizes.
-III-D Dataset Update and Classification
-The dataset is updated with the synthetic data points to form $X_{augmented}\leftarrow X\cup\{x_{synth}\}$, providing an enriched set of training data for machine learning models.
-Leveraging machine learning models on the updated dataset enables the classification tasks for identifying patterns and making predictions.
-In summary, AGCL tackles the challenges of small medical datasets by combining feature extraction, iterative K-means clustering, and synthetic data generation.
-The augmented dataset, comprising both original and synthetic data points, is utilized to train machine learning models, enhancing their performance in classification tasks.
-Ultimately, the algorithm is presented generically in Algorithm 1.
-Input : Dataset $X$, Labels $y$, Parameters $\textit{thrsh},\alpha$
-Output : Augmented dataset $X_{augmented}$
-$X\leftarrow\text{FeatureExtractor}(X)$
-Fit K-means on $X$ to obtain Clusters $C=\{c_{1},c_{2},...,c_{k}\}$ with centers $\{\mu_{1},\mu_{2},...,\mu_{k}\}$
-$X_{synth}\leftarrow\emptyset$
-for $j=1$ to $k$ do
-       $X_{j}\leftarrow\{x\mid x\in X\text{ and }x\in c_{j}\}$
-       $y_{j}\leftarrow\{y\mid x\in c_{j}\}$
-       $n_{c_{j}}\leftarrow\text{size}(X_{j})$
-      if $\text{Unique}(y_{j})$ is False  then
-             Separation $\textit{Separation Criterion}\leftarrow\frac{\text{CSM}(c_{j})}{\text{H}(c_{j})}$
-[…]
-                   Call the algorithm recursively with $X_{j}$, $y_{j}$
-      else
-$r_{j}\leftarrow\begin{cases}0.01\times\frac{1}{d}\sum_{i=1}^{d}\sigma_{j}&\text{if }n_{c_{j}}=1\\ \max(D_{x,\mu_{j}})+0.1\times(\max(D_{x,\mu_{j}})-\min(D_{x,\mu_{j}}))&\text{if }n_{c_{j}}>1\end{cases}$
-             $q\leftarrow\alpha\cdot n_{c_{j}}$
-             $\Sigma_{j}\leftarrow\text{diag}(r_{j}^{2})$
-            for $i=1$ to $q$ do
-                   $x_{new}\sim N(\mu_{j},\Sigma_{j})$
-                   $X_{synth}\leftarrow X_{synth}\cup\{x_{new}\}$
-$X_{augmented}\leftarrow X\cup X_{synth}$
-Algorithm 1 AGCL Pseudocode
+III-D Mapping Synthetic Data to Original Space
+The synthetic data generated in the feature space are mapped back to the original space using a reconstruction network.
+This network is designed to interpret the abstract features of the synthetic data and transform them into a format that closely resembles the original dataset.
+The reconstruction process ensures that the synthetic data retain the structural and contextual characteristics of the real data, making it suitable for further analysis and training.
+By mapping the synthetic data to the original space, the model can leverage the augmented dataset to enhance its learning capabilities while maintaining the real-world properties of the data.
+III-E CNN Model Training and Fine-Tuning
+The Convolutional Neural Network (CNN) is initially trained on the synthetic dataset, which includes only the reconstructed synthetic data.
+This training phase allows the model to learn robust features from the enriched dataset, improving its ability to generalize across diverse data samples.
+Once the initial training is complete, the CNN is fine-tuned on the original dataset to adapt to the specific nuances and patterns present in the real data.
+Fine-tuning ensures that the model retains its generalization capabilities while optimizing its performance for the target task.
+Finally, the trained and fine-tuned CNN performs feature extraction and final classification, using the learned features to accurately identify patterns and make predictions.
+This integrated approach enhances the model’s performance, particularly in scenarios with limited data, by combining the benefits of synthetic data augmentation and targeted fine-tuning.
 IV Dataset
@@
 Examples of these facial expressions are illustrated in Figure 2.
-Not all participants completed all five facial expressions; therefore, the distribution of videos for each emotion is presented in Table I.
 Since each video represents a transition from a neutral face to an extreme emotion, the peak frame of each video is selected for analysis.
-Figure 2: Facial expressions from left to right: anger, happy, disgust, fear and surprise
-TABLE I: Dataset Characteristics
-Emotion Rage Happy Disgust Fear Surprise
-Parkinson’s (PD) 26 25 26 25 29
-Control 28 30 27 30 26
-TABLE II: Associated AUs to each facial expression
+Figure 2: Facial expressions from left to right: anger, happiness, disgust, fear and surprise
+V Experiments
+In this section, we evaluate the effectiveness of the proposed FICAug framework across multiple experimental settings.
+We compare FICAug with baseline models and existing augmentation strategies in both feature space and image space, highlighting its capability to generate informative synthetic samples and improve feature learning.
+V-A Feature Extraction and Feature Space
+TABLE I: Associated AUs to each facial expression
 Emotion Facial Action Unit
@@
 AU15 (Lip Corner Depressor)
-V Experiments
-In this section, we evaluate the AGCL method in various scenarios. Our objective is to provide a thorough assessment of AGCL’s performance under diverse conditions.
-V-A Feature Extraction and Latent Space
 Utilizing Openface [32], we extracted action unit (AU) values for each participant’s picture.
@@
 In this paper, each facial expression is associated with three Aus by analyzing Aus’ mean, concept of each AU associated with facial expression and Pearson Correlation.
-Table II shows the AUs associated with facial expressions.
-TABLE III: Comparison between features of real and synthetic data based on t-test, Levene-test and ks-test
-t-test (p-value) Levene-test (p-value) ks-test (p-value)
+Table I shows the AUs associated with facial expressions.
+V-B Data Splitting for Training, Validation, and Testing
+We employed a Two-Leave-Out cross-validation approach, creating 25 unique folds where each iteration used one control and one patient sample for validation while training on the remainder.
+This ensures comprehensive, unbiased evaluation as all samples serve in validation exactly once.
+The data partitioning maintains balanced age and gender distributions across sets to minimize bias.
+Final predictions combine results from all 25 models through weighted majority voting, where each model’s contribution reflects its validation performance.
+TABLE II: Distribution of training, validation and test sets for each emotion
+Dataset Group Anger Happiness Disgust Fear Surprise
+Train/Val Parkinson 22 18 23 22 22
+Control 19 19 17 19 17
+Test Parkinson 4 7 3 2 7
+Control 9 11 11 11 9
+V-C Statistical evaluations
+TABLE III: Statistical comparison between features of real and synthetic data using t-test, Levene’s test, and Kolmogorov–Smirnov (KS) test, with corresponding p-values.
+t-test Levene-test ks-test
 Emotion Feature Control Parkinson Control Parkinson Control Parkinson
@@
 AU20 0.36 0.79 0.71 0.34 0.57 0.57
-Happy AU15 0.21 0.64 0.23 0.72 0.01 0.98
+Happiness AU15 0.21 0.64 0.23 0.72 0.01 0.98
 AU17 0.01 0.53 0.01 0.94 0.01 0.57
@@
 AU15 0.85 0.15 0.26 0.77 0.17 0.33
-V-B Statistical evaluations
-Three different tests were conducted to verify that the real data and synthetic data do not differ statistically.
+Three different tests were conducted to verify that the real data and synthetic data in feature space do not differ statistically.
 Twenty random samples were selected from each class of the two datasets (real and synthetic).
@@
 Therefore, we can conclude that the two datasets are statistically similar.
-V-C Classification
-TABLE IV: Comparison between accuracy of AGCL, GN, kNNMTD and Baseline on 5 different emotions.
+V-D Evaluation of Feature Space Augmentation Strategies
+The goal of this experiment is to evaluate whether the proposed augmentation strategy in the feature space is effective enough to serve as a reliable foundation for our next steps, including image generation and training CNNs for feature learning.
+We aim to determine whether FICAug can produce meaningful and diverse synthetic samples that perform competitively in classification tasks, making it a valid alternative to other augmentation methods.
+To this end, we compared FICAug against three approaches: a baseline (no augmentation), Gaussian noise (GN) perturbation, and the $k$NNMTD algorithm [19], using four classifiers across five emotions.
+Gaussian noise augmentation involved adding random noise drawn from a normal distribution with varying standard deviations.
+For FICAug, synthetic data points were generated within well-separated clusters using a normal distribution guided by cluster-specific parameters, with grid search used to tune the $\alpha$ value controlling augmentation size.
+As shown in Table IV, the results across validation and test accuracy demonstrate that FICAug generally performs on par with GN and $k$NNMTD, occasionally outperforming them in specific emotions.
+In Rage, FICAug reached a notable test accuracy of 82.82% with MLP.
+The relatively high standard deviation in the validation results is due to the two-leave-out cross-validation strategy.
+With only two samples (one per class) used in each validation fold, even a single misclassification can have a substantial impact on the accuracy, resulting in greater variability across folds.
+For Happiness and Disgust, proposed method maintained strong and competitive performance.
+In Happiness, it reached a high validation accuracy of 81.25% with both MLP and Random Forest, and a top test accuracy of 83.33% using SVM.
+In Disgust, the FICAug again performed best in test accuracy (64.28%) despite Gaussian noise occasionally showing slightly better validation results.
+These findings indicate that while some methods show short-term gains, the proposed strategy generalizes better to unseen data.
+In the Fear emotion, although our method achieved the highest validation accuracy (76.00%) with MLP, its final test results were comparable to those of Gaussian noise.
+In Surprise, FICAug achieved 75.00% in test and 82.00% in cross validation, demonstrating its effectiveness in generating high-quality augmented data.
+However, when aggregating predictions across all emotions using majority voting, the FICAug delivered the most consistent and accurate overall results.
+It achieved the highest test accuracy across all classifiers, and attained the top validation score of 84.09% with Random Forest.
+The final test accuracy reached 77.77% (RF) and 83.33% (MLP), showing significant improvement over the baseline.
+While FICAug does not always significantly outperform other methods in classification accuracy in feature space, its purpose is not solely to boost performance at the feature level.
+Instead, the proposed augmentation strategy in feature space is designed as a structured augmentation mechanism that avoids generating synthetic points in mixed or ambiguous regions of the feature space—areas where different classes intersect.
+This selective sampling is crucial for guiding the model to focus on distinct, class-specific regions, enabling more effective downstream learning.
+Unlike generic noise injection or diffusion-based sampling, FICAug incorporates a class-separation-aware strategy, which we believe is key to generating training data that leads to more meaningful representations when mapped back to the image space.
+TABLE IV: Comparison between accuracy of FICAug, GN, kNNMTD and Baseline on 5 different emotions.
 KNN SVM RF MLP
 Emotion Method Validation Test Validation Test Validation Test Validation Test
-Rage Baseline 54.00 ($\pm$39.80) 69.23 52.00 ($\pm$38.68) 53.84 64.00 ($\pm$38.78) 46.15 66.00 ($\pm$33.82) 53.84
-GN 58.00 ($\pm$39.19) 69.23 56.00 ($\pm$38.26) 53.84 68.00 ($\pm$37.09) 46.15 72.00 ($\pm$31.87) 46.15
-kNNMTD 54.00 ($\pm$39.80) 69.23 56.00 ($\pm$38.26) 53.84 64.00 ($\pm$38.78) 53.84 54.00 ($\pm$42.24) 61.53
-AGCL 60.00 ($\pm$40.00) 76.92 58.00 ($\pm$41.67) 69.23 68.00 ($\pm$39.70) 53.84 74.00 ($\pm$32.00) 53.84
-Happy Baseline 70.83 ($\pm$40.61) 61.11 70.83 ($\pm$43.10) 72.22 79.17 ($\pm$35.11) 66.66 75.00 ($\pm$38.19) 66.67
-GN 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$35.11) 72.22 79.17 ($\pm$35.11) 66.66 81.25 ($\pm$31.66) 77.78
-kNNMTD 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$37.96) 72.22 81.25 ($\pm$34.80) 66.68 68.75 ($\pm$42.85) 61.11
-AGCL 70.83 ($\pm$40.61) 61.11 79.17 ($\pm$35.11) 83.33 81.25 ($\pm$34.80) 72.22 81.25 ($\pm$31.66) 83.33
-Disgust Baseline 68.00 ($\pm$37.09) 46.15 76.00 ($\pm$32.00) 46.15 78.00 ($\pm$28.57) 61.54 70.00 ($\pm$40.00) 69.23
-GN 76.00 ($\pm$34.99) 53.85 80.00 ($\pm$31.62) 38.46 82.00 ($\pm$24.00) 69.23 82.00 ($\pm$27.86) 30.77
-kNNMTD 68.00 ($\pm$37.09) 46.15 76.00 ($\pm$34.99) 46.15 84.00 ($\pm$23.32) 61.54 72.00 ($\pm$37.63) 61.54
-AGCL 72.00 ($\pm$37.63) 69.23 82.00 ($\pm$31.24) 53.84 84.00 ($\pm$30.72) 69.23 84.00 ($\pm$27.28) 76.92
-Fear Baseline 60.00 ($\pm$40.00) 69.23 68.00 ($\pm$37.09) 84.62 62.00 ($\pm$38.16) 61.53 62.00 ($\pm$38.16) 53.84
-GN 64.00 ($\pm$36.11) 69.23 66.00 ($\pm$39.29) 76.92 66.00 ($\pm$36.66) 61.53 70.00 ($\pm$37.42) 76.92
-kNNMTD 60.00 ($\pm$40.00) 69.23 68.00 ($\pm$39.70) 84.62 72.00 ($\pm$31.87) 53.84 68.00 ($\pm$39.70) 61.53
-AGCL 62.00 ($\pm$38.16) 76.92 70.00 ($\pm$40.00) 84.62 74.00 ($\pm$37.74) 69.23 72.00 ($\pm$37.63) 92.30
+Rage Baseline $54.00(\pm 39.80)$ $69.23$ $52.00(\pm 38.68)$ $53.84$ $64.00(\pm 38.78)$ $46.15$ $54.00(\pm 42.24)$ $76.92$
+GN $58.00(\pm 39.19)$ $69.23$ $56.00(\pm 38.26)$ $53.84$ $64.00(\pm 37.09)$ $46.15$ $64.00(\pm 32.00)$ $76.92$
+kNNMTD $54.00(\pm 39.80)$ $69.23$ $56.00(\pm 38.26)$ $53.84$ $64.00(\pm 38.78)$ $53.84$ $54.00(\pm 39.80)$ $69.23$
+FICAug $60.00(\pm 40.00)$ $76.92$ $58.00(\pm 41.67)$ $69.23$ $68.00(\pm 39.70)$ $53.84$ $64.00(\pm 38.78)$ $84.61$
+Happiness Baseline $70.83(\pm 40.61)$ $61.11$ $70.83(\pm 43.10)$ $72.22$ $79.17(\pm 35.11)$ $66.66$ $75.00(\pm 38.19)$ $77.77$
+GN $70.83(\pm 40.61)$ $61.11$ $70.83(\pm 35.11)$ $72.22$ $79.17(\pm 35.11)$ $66.66$ $79.17(\pm 35.11)$ $66.66$
+kNNMTD $70.83(\pm 40.61)$ $61.11$ $70.83(\pm 37.96)$ $72.22$ $79.17(\pm 34.80)$ $66.66$ $68.75(\pm 42.85)$ $72.22$
+FICAug $70.83(\pm 40.61)$ $61.11$ $79.17(\pm 35.11)$ $83.33$ $81.25(\pm 34.80)$ $72.22$ $81.25(\pm 34.80)$ $77.77$
+Disgust Baseline $68.00(\pm 37.09)$ $42.85$ $76.00(\pm 32.00)$ $42.85$ $78.00(\pm 28.57)$ $57.14$ $70.00(\pm 40.00)$ $57.14$
+GN $76.00(\pm 34.99)$ $50.00$ $80.00(\pm 31.62)$ $35.71$ $82.00(\pm 24.00)$ $64.28$ $80.00(\pm 31.62)$ $35.71$
+kNNMTD $68.00(\pm 37.09)$ $42.85$ $76.00(\pm 34.99)$ $42.85$ $84.00(\pm 23.32)$ $57.14$ $68.00(\pm 42.14)$ $50.0$
+FICAug $72.00(\pm 37.63)$ $64.28$ $82.00(\pm 31.24)$ $50.00$ $84.00(\pm 30.72)$ $64.28$ $82.00(\pm 27.86)$ $64.28$
+Fear Baseline $60.00(\pm 40.00)$ $69.23$ $68.00(\pm 37.09)$ $84.62$ $58.00(\pm 39.19)$ $69.23$ $62.00(\pm 40.69)$ $61.53$
+GN $64.00(\pm 41.28)$ $69.23$ $68.00(\pm 39.70)$ $76.92$ $64.00(\pm 43.63)$ $69.23$ $70.00(\pm 37.42)$ $69.23$
+kNNMTD $60.00(\pm 40.00)$ $69.23$ $68.00(\pm 39.70)$ $76.92$ $72.00(\pm 31.87)$ $53.84$ $72.00(\pm 37.63)$ $61.53$
+FICAug $62.00(\pm 43.08)$ $69.23$ $72.00(\pm 37.63)$ $76.92$ $74.00(\pm 34.99)$ $69.23$ $76.00(\pm 34.99)$ $69.23$
 Surprise Baseline $68.00(\pm 39.70)$ $56.25$ $78.00(\pm 34.87)$ $43.75$ $70.00(\pm 34.64)$ $43.75$ $72.00(\pm 34.87)$ $62.50$
-GN 72.00 ($\pm$37.63) 50.00 70.00 ($\pm$37.42) 43.75 80.00 ($\pm$31.62) 50.00 76.00 ($\pm$34.99) 43.75
-kNNMTD 68.00 ($\pm$39.70) 56.25 68.00 ($\pm$39.70) 50.00 68.00 ($\pm$34.29) 50.00 72.00 ($\pm$37.63) 56.25
-AGCL 70.00 ($\pm$37.42) 62.50 78.00 ($\pm$34.87) 50.00 82.00 ($\pm$31.24) 56.25 84.00 ($\pm$30.72) 75.00
-Voting Baseline 68.18 61.11 75.00 66.66 77.27 61.11 65.90 72.22
-GN 70.45 66.66 70.45 61.11 81.81 66.66 77.27 66.66
-kNNMTD 68.18 61.11 77.27 66.66 79.54 66.66 77.27 66.66
-AGCL 75.00 77.77 79.54 77.77 90.90 77.77 84.09 83.33
-In the previous section, we analyzed the data statistically.
-In this section, we evaluate how the data behaves when subjected to machine learning algorithms for classifying Parkinson’s and control data.
-A two-leave-out cross-validation method was applied to the training data, resulting in 25 unique folds.
-Each fold served as an independent subset for training separate models using the same hyperparameters.
-Throughout the cross-validation process, hyperparameter tuning was carried out to identify the optimal hyperparameters based on performance across all folds.
-The test data consisted of 11 control samples and 7 Parkinson’s disease samples.
-Instead of selecting a single model, all 25 models generated during the cross-validation process were applied to the test data.
-Their predictions were aggregated using a weighted majority voting approach to produce the final classification.
-By comparing the effectiveness of AGCL against baseline, perturbation mechanisms and kNNMTD [16], we aim to evaluate its performance in classifying facial expressions.
-The baseline approach serves as a reference point without any augmentation, providing a foundational measure of classification performance.
-In contrast, the perturbation-based synthetic data generation mechanism is stated as: $x^{s}=x_{i}+\epsilon$ where $\epsilon$ is a noise sampled from a certain distribution.
-According to Gaussian approach, $\epsilon$ is sampled with $N(0,\sigma)$.
-This method usually utilizes n-way marginals for purposes of creating variability in the categorical features [31].
-Therefore, we employed the Gaussian approach to add noise to each sample in the latent space.
-To elucidate the impact of different levels of noise, we consider a range of standard deviations ($\sigma$= 0.001, 0.005, 0.01, 0.025, 0.05, 0.075, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.8, 1]) to identify the optimal performance of Gaussian Noise (GN) for comparison with the AGCL method.
-Moreover, We estimated the number of generated data points in the cluster for AGCL through grid search, gradually testing the $\alpha$ value from 2 to 60 with a step size of 5.
-Table IV shows the performance of algorithms across four machine learning classifiers in screening Parkinson across five emotions: Rage, Happy, Disgust, Fear, and Surprise.
-To aggregate results obtained from five facial expressions, we employ the majority voting integration method, by voting from all facial expressions for each individual.
-This ensemble method helps in reducing the possibility of bias or errors that come with separate facial expressions, making it a more accurate and diverse approach.
-The metrics reported include both cross-validation and test accuracy.
-For the rage emotion, AGCL demonstrated superior performance compared to other methods across all classifiers.
-The best validation accuracy was achieved using the MLP classifier, where AGCL reached 74.00%, whereas the baseline validation accuracy was lower at 66.00%.
-In the test results, AGCL also led with the highest accuracy of 76.92% with kNN, outperforming the other algorithms test accuracy of 69.23%.
-In the happy emotion, AGCL achieved the highest validation accuracy of 81.25% with MLP, significantly improving over the kNNMTD’s validation accuracy of 68.75%.
-The test accuracy of 83.33% with both SVM and MLP shows the effectiveness of AGCL in generating high-quality synthetic data that enhances classification performance, compared to the GN and kNNMTD.
-While AGCL performed well in test accuracy compared to the baseline and other augmentation methods, the overall improvement in cross-validation was less pronounced than in other emotions.
-The validation accuracy of 81.25% with RF and MLP showed moderate gains, which suggests that the happy expression might be easier for classifiers to distinguish even without extensive data augmentation.
-For the disgust emotion, AGCL also led to noticeable performance improvements.
-The best validation accuracy of 84.00% was achieved using the Random Forest classifier.
-AGCL also reached a test accuracy of 76.92% with MLP, which is the best test accuracy across all four algorithms.
-The fear emotion further highlights the advantages of AGCL in small dataset classification.
-AGCL achieved the highest validation accuracy of 74.00% with the Random Forest classifier, while the GN struggled with only 66.00%.
-On the test set, AGCL achieved its best accuracy with MLP, reaching an outstanding 92.30%, far exceeding the other algorithms.
-In the surprise emotion, AGCL continued to outperform the baseline and other methods.
-The highest validation accuracy was 84.00% with MLP for AGCL.
-In the test phase, AGCL again demonstrated strong performance, achieving 75.00% accuracy with MLP, surpassing the baseline’s test accuracy of 62.50%.
-When results across all emotions were aggregated using majority voting, AGCL achieved the highest overall test accuracy of 83.33% with the MLP classifier.
-This marks a significant improvement over the other algorithms.
-The validation accuracy of 90.90% with Random Forest and 84.09% with MLP further confirms AGCL’s effectiveness in augmenting small medical datasets and improving classification results.
+GN $68.00(\pm 37.63)$ $50.00$ $70.00(\pm 37.42)$ $43.75$ $80.00(\pm 31.62)$ $50.00$ $78.00(\pm 31.87)$ $43.75$
+kNNMTD $68.00(\pm 39.70)$ $56.25$ $68.00(\pm 39.70)$ $50.00$ $68.00(\pm 34.29)$ $50.00$ $74.00(\pm 34.99)$ $62.50$
+FICAug $70.00(\pm 37.42)$ $62.50$ $78.00(\pm 34.87)$ $50.00$ $82.00(\pm 31.24)$ $56.25$ $82.00(\pm 31.24)$ $75.00$
+All Emotions Baseline $68.18$ $61.11$ $72.72$ $66.66$ $75.00$ $66.66$ $68.18$ $72.22$
+GN $70.45$ $66.66$ $72.72$ $61.11$ $82.36$ $66.66$ $81.81$ $66.66$
+kNNMTD $68.18$ $61.11$ $77.27$ $66.66$ $79.54$ $66.66$ $75.00$ $77.77$
+FICAug $75.00$ $72.22$ $79.54$ $72.22$ $84.09$ $77.77$ $81.81$ $83.33$
+V-E GANimation and Synthetic Face Generation
+To initiate the generation process in the original image space, we first utilized the Generated Photos tool [36] to create a collection of neutral facial images.
+This tool employs GAN-based models to synthesize realistic human faces.
+To increase the realism and relevance of the dataset, we specifically selected images of elderly individuals with neutral expressions (i.e., without any emotional cues).
+These images provided a strong foundation for applying the synthetic AUs.
+Examples of these synthesized neutral faces are shown in Figure 3.
+Figure 3: AI-generated neutral face images used for applying Action Units
+The GANimation  [ganimation] model is a generative adversarial network specifically designed for facial expression synthesis.
+It allows for precise and continuous control over facial movements by applying AUs to a given face, without requiring paired training data.
+Unlike traditional conditional GANs, GANimation uses attention and deformation mechanisms to focus only on the facial regions that need to be modified, leading to more natural and expressive outputs.
+Its architecture is particularly well-suited for generating identity-preserving facial expressions across various emotional states.
+To improve the performance and accuracy of the GANimation model, we fine-tuned it using the generated neutral images.
+The model was fine-tuned for 20 epochs on this dataset, allowing it to adapt to the specific characteristics of our data and learn relevant facial features.
+In the next step, we fed the GANimation model with the synthesized neutral images along with facial AUs generated for each class using the clustering algorithm.
+These AUs were created in two ways: (1) a set of dominant AUs selected for each emotion based on our proposed method involving clustering, cluster evaluation, and re-clustering, and (2) additional AUs sampled from a normal distribution to increase diversity and realism.
+The normal distribution helped expand the variation in generated expressions and improve the representation of emotional diversity.
+Finally, GANimation applied the given AUs to the neutral faces and produced diverse synthetic facial expressions for each class in the dataset.
+The neutral faces used for both control and Parkinson’s groups were identical; the only difference lay in the AUs applied to them.
+Examples of the generated images for each emotion and both groups are illustrated in Figures 4 and 5.
+Figure 4: Synthetic facial expressions generated for control individuals using GANimation
+Figure 5: Synthetic facial expressions generated for Parkinson’s patients using GANimation
+To evaluate the quality of the generated images, we used the OpenFace toolkit to extract AUs from the synthetic images.
+These extracted AU values were then compared to the input AUs fed to the GANimation model, and the Mean Absolute Error(MAE) was calculated.
+Table V presents the MAE values for each emotion and both subject groups.
+The low error values confirm the model’s ability to accurately reproduce the intended AUs.
+TABLE V: Evaluation of GANimation using MAE
+Emotion Rage Happiness Disgust Fear Surprise
+Parkinson’s 0.17 0.09 0.08 0.19 0.09
+Control 0.16 0.05 0.07 0.06 0.07
+V-F Training the Convolutional Neural Network Model
+To train the proposed model, we first used ResNet18 and trained it on the synthetic emotional face images.
+Since the neutral faces were identical for both groups, this initial training phase encouraged the model to focus on differences in emotional expression and Action Unit patterns, rather than on irrelevant visual differences between individual faces.
+After this pretraining phase, the model was fine-tuned on the original dataset.
+Throughout the training process, various data augmentation techniques were employed to improve the model’s performance.
+During the first phase, when training on synthetic images, transformations such as horizontal flipping, rotation, and perspective distortion were applied.
+In the second phase, where the model was fine-tuned on the original dataset, Gaussian noise was additionally introduced.
+These augmentation strategies helped increase the diversity of training samples and enhanced the model’s generalization capability.
+Table VI presents the classification results of FICAug compared to three baselines: a standard ResNet18, a DINO-pretrained ResNet50, and the Masked Siamese Networks(MSN) model using facebook/vit-msn-small.
+All models were trained under identical settings, with consistent data splits, augmentation techniques, training epochs, and optimizer configurations.
+This ensures a fair comparison between the methods.
+Across all five emotions as well as the aggregated All Emotions category, FICAug consistently outperforms the baselines in both validation and test accuracy.
+The relatively high standard deviation observed in validation results stems from the use of a two-leave-out cross-validation strategy.
+Since each fold contains only two samples for validation (one from each class), any misclassification significantly affects the accuracy, leading to larger variance across folds.
+TABLE VI: Comparison of classification accuracy between FICAug and other algorithms
+Emotion Method Cross Validation Test
+Rage Baseline $62.00(\pm 38.15)$ $53.84$
+DINO $44.00(\pm 29.00)$ $30.76$
+MSN $70.0(\pm 37.41)$ $61.53$
+Proposed $74.00(\pm 37.73)$ $69.23$
+Happiness Baseline $72.91(\pm 35.29)$ $72.22$
+DINO $60.41(\pm 40.77)$ $55.55$
+MSN $83.33(\pm 27.63)$ $55.55$
+FICAug $81.25(\pm 34.80)$ $77.77$
+Disgust Baseline $82.00(\pm 27.27)$ $71.42$
+DINO $62.00(\pm 35.44)$ $50.00$
+MSN $82.00(\pm 27.85)$ $64.28$
+FICAug $86.00(\pm 26.53)$ $78.57$
+Fear Baseline $70.00(\pm 34.64)$ $69.23$
+DINO $52.00(\pm 29.93)$ $30.76$
+MSN $76.00(\pm 32.00)$ $69.23$
+FICAug $78.00(\pm 34.87)$ $76.92$
+Surprise Baseline $68.00(\pm 37.09)$ $75.00$
+DINO $40.00(\pm 31.62)$ $56.25$
+MSN $80.00(\pm 34.64)$ $62.50$
+FICAug $68.00(\pm 37.09)$ $75.00$
+All Emotions Baseline $70.96$ $72.22$
+DINO $50.00$ $66.66$
+MSN $68.18$ $72.22$
+FICAug $88.63$ $94.00$
+In the Rage emotion, FICAug achieves a test accuracy of 69.23%, surpassing the baseline ResNet18 (53.84%), DINO (30.76%), and MSN (61.53%).
+For Happiness, although the MSN model showed the highest cross-validation accuracy (83.33%), its test performance dropped significantly to 55.55%.
+FICAug, in contrast, maintained strong generalization with a test accuracy of 77.77%.
+In the case of Disgust, FICAug achieved the highest cross-validation accuracy (86.00%) and test accuracy (78.57%), outperforming all three baselines.
+Moreover, for the Fear category, our method again demonstrated superior performance, achieving the best test accuracy of 76.92%, compared to 69.23% for both the baseline and MSN, and only 30.76% for DINO.
+In the Surprise category, FICAug matched the baseline with a test accuracy of 75.00%, while both DINO and MSN underperformed at 56.25% and 62.50%, respectively.
+Aggregating results across all emotions, FICAug achieved a substantial improvement, reaching a test accuracy of 94.00%, compared to 72.22% for both the baseline and MSN, and 66.66% for DINO.
+The corresponding validation accuracy was also the highest at 88.63%, further confirming the generalization capabilities of our approach.
+These results clearly demonstrate that FICAug, which combines feature-space clustering and synthetic image generation, is an effective strategy for enriching training data and improving classification performance—particularly in small data regimes.
+Unlike self-supervised approaches, our method generates new training samples by mapping existing, clustered latent features back to the image space, guided by an augmentation strategy designed to increase diversity and coverage of expression patterns.
+By training a CNN on these generated images, the model is not simply learning from duplicated or noised data—it is exposed to novel expressions synthesized from meaningful variations in the original features.
+This enables the extraction of new, discriminative representations that better reflect subtle emotional cues, as evidenced by the method’s superior performance.
+V-G Comparison of Feature Quality: Feature Space vs. Image-Based Learning
+In the final analysis, we compare classification performance between two strategies: direct classification using augmented latent features versus training a CNN on synthetic facial images reconstructed from those features.
+As shown in Figure 6, learning from image-based data using CNN leads to a substantial performance boost.
+The validation accuracy increased to 88.63%, compared to 81.81% in the feature space approach.
+Similarly, test accuracy improved markedly, reaching 94.00%, as opposed to 83.33% with feature space classification.
+This improvement highlights the strength of our approach: by generating synthetic images from existing latent features and training a CNN on them, the model is able to extract new, richer representations—effectively learning new features from data derived from itself.
+This self-guided learning process enables the network to go beyond the original feature space and develop more expressive and discriminative patterns, ultimately improving accuracy.
+Figure 6: Comparison of validation and test accuracy between latent space classification and image-based CNN training across all emotions.
 VI Conclusion
-In this paper, we introduced the Artificial Data Point Generation in Clustered Latent Space (AGCL) method to address the challenges of limited training data in medical applications, particularly in the classification of Parkinson’s disease through facial expression analysis.
-By generating synthetic data points in a clustered latent space, AGCL enhances the performance of machine learning models in scenarios with small datasets.
-The framework employs feature extraction, K-means clustering, and a separation criterion to guide synthetic data generation, ensuring that only well-separated clusters contribute to data augmentation.
-Experimental results across multiple classifiers and emotions demonstrate that AGCL consistently outperforms baseline methods, as well as other data augmentation techniques like Gaussian noise and kNNMTD.
-AGCL achieved the highest overall test accuracy of 83.33% through majority voting, confirming its effectiveness in improving classification in small datasets.
-AGCL’s ability to generate synthetic data points tailored to the structure of the dataset offers significant advantages in enhancing the generalization capabilities of machine learning models.
-This method holds promise for broader applications in medical domains and other fields where data scarcity is a critical issue.
-Future work may explore the extension of AGCL to different datasets and further refinement of clustering techniques to increase its applicability across a wider range of tasks.
+This study presented FICAug, a feature-to-image data augmentation framework aimed at improving machine learning performance under small-data conditions.
+The method first generates synthetic samples in the feature space using k-means clustering and Gaussian sampling.
+To avoid increasing data complexity, data are not generated in regions where class overlap exists.
+This targeted augmentation improves diversity while maintaining the structural integrity of the original distribution.
+Experimental evaluations on a facial expression dataset for Parkinson’s disease screening demonstrated that the generated data preserved key statistical properties of the original samples and significantly improved classification accuracy across multiple classifiers.
+To enhance representation learning, the synthetic features were mapped back into the image space using a GANimation-based generative model.
+Notably, FICAug achieved a cross-validation accuracy of 88.63%, significantly surpassing the best baseline performance of 70.96%.
+This highlights the ability of FICAug to extract richer and more discriminative features from structured synthetic data.